OpenStudy (anonymous):

"Find the equation of the line MN that has the given information:" m= -2, M (4,5)

OpenStudy (anonymous):

ok so it passes through point (4,5) and the slope is -2 y - y1 = m (x-x1) y - 5 = -2 (x-4) This is point slope form what form do you need it in. you may be able to just leave it like this...

OpenStudy (anonymous):

@Amineh0923 ?? you there bruh lol

OpenStudy (anonymous):

lol ya im here just trying to make sense of it

OpenStudy (anonymous):

ok do I need to explain better?

OpenStudy (anonymous):

yes if that's ok

OpenStudy (anonymous):

ok so when you are given slope and a point on the graph, the first thing you need to do is put it in point slope form: y - y1 = m (x-x1) y1 = the y-value of the coordinate you are given x1 = the x-value of the coordinate you are given m = the slope In this problem y1 = 5 x1 = 4 m = -2 Just plug it into the formula Can you do that part

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes a lot better thank you

OpenStudy (anonymous):

Awesome!